OpenStudy (anonymous):

The distance a car travels varies directly with the number of hours driven. If a car travels 210 miles in 3 hours, write an equation to represent this direct variation

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

210=70(3)+b can you solve for b?

OpenStudy (anonymous):

this i have more can u help

OpenStudy (anonymous):

i could try lets try finishing this one first,, so solve for b

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

simplify 70*3 then subtract from both sides to solve for b...

OpenStudy (anonymous):

do it x the numbers

OpenStudy (anonymous):

yes multiply 70 and 3

OpenStudy (anonymous):

is 210

OpenStudy (anonymous):

yes so now subtract 210 from both sides of the equal sign

OpenStudy (anonymous):

what does that look like the equal sign

OpenStudy (anonymous):

70-210 is -140

OpenStudy (anonymous):

and 3 - 210 is -207

OpenStudy (anonymous):

no 210-210=210-210+b so b=0 so our equation is y=70x

OpenStudy (anonymous):

omg thank you

OpenStudy (anonymous):

do u have more time ?????

OpenStudy (anonymous):

no problem, yea post a new thread and tagg me in it

OpenStudy (anonymous):

ok